0.1 mole of a compound of S and cl on oxidation with concentration of HNO3 and subsiquent treatment gave 0.1 mole of PbCl2 and 0.2 mole of BaSo4 as precipitates.Determine the molecular formula of the compound.

Grade:12

1 Answers

askiitian.expert- chandra sekhar
14 years ago

Hi pallavi,

Let the compound be sxcly

no of moles of Sxcly=0.1

therefore,S atoms=0.1×X

              cl atoms=o.1×y

no of moles of Pbcl2 =o.1

therefore cl atoms=0.1×2=0.2

no of moles of BaSo4=0.2

therefore S atoms=o.2 moles


0.1× x=0.2


→x=2

→0.1×y=0.2

   y=2

therefore the molecular formula of the compound is S2Cl2.